MIMD-Style Parallel Programming Based on Continuation-Passing Threads

Computation Structures Group Memo 355 March 7, 1994

Michael Halbherr, Yuli Zhou and Chris F. Joerg

This report describes research done at the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-92-J-1310.